As a salesperson at Russell's Furniture, Abdul has a choice between two options for receiving monthly pay. With one option, he receives a base salary of $1132 and gets an additional 4% commission on the sales he makes. With the other option, he receives a base salary of $300 and gets an additional 12% commission on the sales he makes. What amount of sales must he have for the pay from the two options to be equal?








Answer by mananth(16949)

Let the sales be x
Option A = Option B
1132 +0.04x=300+0.12x
1132-300=0.08x
832=0.08x
/0.08
832/0.08=x
x=10,400 sales